# Modeling wildfire dynamics through a physics-based approach incorporating fuel moisture and landscape heterogeneity

### v1

@@ 3.2 Time stepping
We use the Strong Stability Preserving Runge–Kutta 3 (SSP-RK3) scheme [25, 49] for stepping in time the semi-discrete Eq. (67) as follows
@@ 4 Numerical results
We start with simulations in one spatial dimension and study the influence of different parameters. For the 1D simulations, we consider a computational domain $[0,500]$ m and set as initial condition
@@
For all cases, we use a computational mesh composed of $N_{x}=1600$ cells and $\mathsf{CFL}=0.1$. Default parameters from Tabs. 24 are used, unless other configuration is specified.
@@ 4.1 Comparison of combustion functions
We compare two approaches for the combustion function presented in Sec. 2.5. Using the fixed parameters in Tab. 4, we determine the parameter $A_{c}$ of the point-wise constant combustion function in Eq. (37) such that it approximates the nonlinear Arrhenius law Eq. (35) in the relevant temperature domain. Fig. 3(a) shows this comparison. The relevant temperature domain is $[T_{pc},T_{\mathrm{up}}]$, where $T_{\mathrm{up}}$ approximates an upper bound for temperatures in a fire.
@@
Figure 3: Comparison of the two combustion functions Eqs. (35) and (37), and the solutions of Eq. (27) using them.
For comparing the influence of the combustion function, we set the advection term zero ($w=0$) and compute the solutions of the ADR-model for a moisture level $M=0.1$. The direct comparison of the solutions in Fig. 3(b) shows that both, the traveling wave profile and the traveling wave speed are almost identical for the two combustion functions. The maximal temperature is slightly higher for the combustion function using the Arrhenius law. This difference has its origin in the approximation in Fig. 3(a), where the constant combustion function has a lower value for large temperatures than the Arrhenius combustion function. The differences of the solutions can be reduced even further by applying calibration techniques for the parameter $A_{c}$. Because of the good approximation, the assumption of a constant reaction rate instead of the Arrhenius rate is convincing.
@@
In the further investigations, we therefore use the piecewise constant reaction rate in Eq. (37) instead of the nonlinear Arrhenius law.
@@ 4.2 Influence of radiation models
The radiation leads to a nonlinear diffusion term, compare Eqs. (17) and (23). When studying the influence of various mechanisms on the system’s behavior in [42], the nonlinear diffusion by radiation was not considered, and the diffusion was modeled only by conduction leading to a linear diffusion term with a modified larger constant $k_{c}$. Here, we first investigate the differences in solutions caused by the simplification of a purely linear diffusion compared to the combined effect of radiation and conduction.
Figure 4: Comparison of the solution profiles with and without radiation. The parameter $k_{c}$ differs in the two simulations for compensating the lack of radiation.
@@ 4.3 Influence of fuel moisture models
In Sec. 2.6, we present different models for the fuel moisture and the effect on the specific heat. Here, we compare the different models for some scenarios.
@@ 4.3.1 Comparison of three moisture models
First, we consider a green wood to total wood ratio $r=1$, meaning that all wood is green wood, and a moisture content of $M=0.1$ under no wind conditions. The rest of the parameters are set as default. In this case, the only differences between the simple and the complete two-stage moisture models S2M and C2M is that the latter considers a correction factor, $A_{i}$, in the specific heat approximation Eq. (61) below wood fibre saturation (moisture content below 0.3). Note that the value of $M=0.1$ used in this test case would be unrealistic for green wood, but it has been chosen as an example for comparison of the models.
@@
Figure 5: Comparison of the two-stage moisture models (S2M, C2M) with the three-stage moisture model S3M. The moisture content is $M=0.1$, the wind speed is $w=0$, and the live/dead ratio is $r=1$.
@@
For a further investigation of the traveling wave profile, we regard the phase portraits of the solutions. This method highlights differences in the solution behavior very well, compare the use in [42]. We fix a certain point $x$ where the traveling wave will pass by after some time. This point $x$ should not be too close to the maximum of the initial data for ensuring that the traveling wave front is already in its stable shape. For this point $x$, we plot the solution values $T(x,t)$ and $Y(x,t)$ over each other, leading to the phase plot in Fig. 6(left). The phase plot shows that the traveling wave front is following nearly the same $T$-$Y$-curve, independent of whether a two-stage or a three-stage model was used. Additionally, we plot the time-dependent solutions for this fixed point $x=350$ m, which shows the profiles in Fig. 6(right). This figure shows again the slightly slower rate of spread of the two-stage models. Note that Fig. 6(right) shows the time-dependency, a faster travelling wave therefore reaches the fixed spatial location therefore in shorter time.
Figure 6: Comparison of two-stage and three-stage moisture models with fixed $x=350$ m, and parameters $M=0.1$, $r=1$, $w=0$. Left: Phase portrait; right: time-dependent solutions.
In the next step, we will better visualize the differences between the 2- and 3-stage fuel moisture models. The 3-stage model explicitly represents the latent heating process into account, resulting in some plateau in Fig. 1(b). This difference is not visible in the phase portrait in Fig. 6, because it takes place while $Y=1$, following a straight line with increasing temperature. In contrast, Fig. 7 shows very well the differences of the time-dependent temperature curves at a fixed $x=350$ m during the heating processes. The parameters are fixed as in the previous figures. The solution using the S3M model shows the latent phase with a slightly increasing plateau, due to the width of the evaporation interval in Eq. (48). This evaporation plateau leads to a slowed heating process for temperatures around the evaporation temperature $T_{w}$. Contrarily, the S2M and C2M models show a monotonically increasing heating process without the evaporation plateau. In these models, the effect of latent heating is treated as a sensible heating and therefore evaporation cannot be explicitly represented.
Figure 7: Heating process for the two-stage models S2M, C2M, and the three-stage model S3M. The evaporation plateau leads to a slowed heating process for temperatures around the evaporation temperature $T_{w}$. Fixed parameters are $x=350$ m, $r=1$, $M=0.1$ and $w=0$.
@@
In the next step, we fix the S2M model and investigate the effect of varying moisture levels.
@@ 4.3.2 Influence of moisture content
For the following investigations, we use the simple two-stage moisture model S2M in Sec. 2.6.2. Further, we do not regard advection effects and set $w=0$. In comparison to the last section, we now vary the moisture content $M$.
Fig. 8 shows the solution at $t=4000$ s for moisture contents $M=0$, $M=0.1$ and $M=0.2$. The traveling wave profiles in Fig. 8 show a decrease in the maximal temperature and in the traveling wave speed with increasing moisture content $M$. The surviving biomass fraction $Y$ is identical for all three values of $M$ even though the maximal temperature is different in all cases. In the complex system behavior, the larger maximal temperature values are compensated by the faster traveling wave speeds. Note that this inverse dependency between maximal temperature and traveling wave speed is due to the nonlinear radiation term. In previous investigations [37], it was shown that a higher rate of spread is connected to a larger remaining biomass fraction $Y$. In this case now, the higher maximal temperature would lead to a smaller remaining biomass fraction, but the faster rate of spread compensates this effect.
Figure 8: Solution profiles with the two-stage moisture model S2M for different moisture levels. Fixed values are $r=1$ and $w=0$.
@@
The preheating time is nonlinear related to the moisture level. In particular, for this parameter setting, a moisture level of $M\geq 0.4$ stops the traveling wave behavior. The longer the preheating time is, the slower are the traveling waves because it needs longer to start the combustion process when the traveling wave front pushes forward.
@@
Figure 9: Comparison of different moisture levels in the S2M model with fixed $x=350$ m, $w=0$, $r=1$. Left: phase portrait; right: time-depending solution profiles.
@@
Figure 10: Time-dependent temperature curves for three neighboring discretization cells (upper), and specific heat values in the three cells (lower). The moisture model is S2M with $M=0.2$, $w=0$, $r=1$ and $x=350$ m.
The occurrence of bumps is independent of the discretization parameters and desired for describing the temperature dependent evaporation process. They do not change the traveling wave behavior, see Fig. 9. Therefore, we notice them as a feature of the model, not distracting any other observations.
@@ 4.4 Investigation of the parameter-dependency of ROS
We use the S2M model for some further investigations on the dependency of the rate of spread on the moisture content, the wind velocity, the relative humidity and the surface coverage. These investigations lead the path towards more realistic scenarios and deepen the understanding of the influence of moisture models.
@@
Fig. 11(a) shows the rate of spread (ROS) of the model in Eq. (27) with respect to the total moisture content, $M$, computed using the model S2M in Eq. (53). In the case without wind, $w=0$, we find a stopping of the traveling wave for larger moisture contents. The fire front stops and the fire is not spreading any further. This solution’s behavior is different for situations with small wind, $w=1$ or $w=2$ m/s. In these cases, the moisture content slows down the rate of spread but does not lead to a stopping of the fire front.
@@
Figure 11: Dependency of the rate of spread (ROS) on the moisture content $M$ and the wind velocity $w$. The S2M model is used for modeling the moisture.
@@
The observation that the traveling wave solution with some wind $w>0$ does not stop is in accordance with practical observations, although the underlying mechanisms differ in reality and in this model. In reality, a strong wind may lead to a heating of the fuel far ahead of the traveling wave front and therefore to an evaporation of the fuel moisture before the fire front arrives. This long-distance effect, however, is not represented in the current mathematical model. Here, the model only considers local radiative effects through the Rosseland approximation, which acts as a diffusion mechanism. As a result, heat is transferred between adjacent areas in a manner similar to conduction. To properly model this effect, we would need to include the non-local effect of radiation as proposed in [23, 52]. This implies the resolution of additional equations for the radiation, thus increasing the complexity of the model.
@@
The impact of increasing fuel moisture on the ROS would be higher if, instead of considering a constant factor $\beta$, we computed it using the expression in Eq. (34) thereby introducing a dependency $\beta=\beta(M)$. As $M$ increases, $\bar{c}_{p,f}$ also increases, leading to a decrease in $\beta$, which results in a larger reduction of the ROS for high moisture and wind values. Nevertheless, given the level of simplification in our model, we limit ourselves to a constant value for $\beta$ and defer this option for future research. This approach could improve the results for strong wind situations, where the effect of moisture is underestimated.
@@
In Sec. 2.3, the connection between wind velocity and topographies with constant slopes was drawn. Here we want to recall the main idea: Even though the investigations in Fig. 11 were carried out for a wind velocity, they are valid as well for a non-flat topography with a constant slope. Further effects of the topography on the spread of fire will be studied with simulation in two spatial dimensions in Sec. 4.5.
@@
In the next step, we vary the relative humidity and the live to dead ratio $r$ of the fuel. The live to dead ratio is relevant only for the complete two-stage moisture model C2M, which we will use for the following investigations.
@@
Figure 12: Dependencies of the rate of spread (ROS) with the C2M moisture model on the relative humidity RH ($\phi$) with varying wind velocity $w$ for $r=0.0$, and varying green wood moisture content $M_{g}$ for $r=0.5$.
@@
The two figures show a behavior that is expected from literature: In [35, Fig. 6], the comparison of different (empirical) models with simulation results is shown. All the models there show a decaying tendency. The different empirical models describe the decay either by a linear or an exponential decay function. Our model results differ from this behavior in the case of very high relative humidity values, where we observe an additional drop. As a relative humidity near 100% is not the typical wildfire condition, we rate the fit of our model to the literature values still to be good.
@@
The dependence of the rate of spread on the bulk density of the fuel has been a widely studied topic for decades [16, 19]. This dependency is highly significant in heterogeneous environments, and recent ADR models have demonstrated the ability to accurately reproduce it [61]. Here, we study the influence of the fuel bulk density, $\bar{\rho}_{f}R_{f,0}$ in Eq. (9), in the ROS. For this numerical experiment, we consider a fixed fuel density (see Tab. 2) and change the fuel volume ratio, $R_{f,0}$, also called packing ratio in this context.
We compare three different models for the dependency of the bulk density on the fuel load. Eqs. (6) and (7) give the first approximation, where both, the gaseous and solid phase, are regarded. Additionally, their portion varies in time due to the combustion process. We name this modeling approach as ’mixture (time varying)’. The next approximation is given by Eq. (8), taking into account the mixture but not the time variation. As the simplest approximation, Eq. (9) does not depend on the gaseous phase but regards only the fuel (solid phase).
Figure 13: Rate of spread (ROS) depending on the bulk density and the packing ratio for the three different models. The simulations are without moisture $M=0$ and wind $w=0$.
Fig. 13 shows the dependency of the rate of spread on the fuel bulk density for the three approaches. The results evidence that there are two well-defined regions. For lower values of the fuel bulk density (and packing ratio), there is a transition from no fire conditions to wave propagation, increasing the ROS up to a maximum value. For larger densities (and packing ratio), the ROS decreases slowly, in accordance with the results reported in the literature. We believe that the role of the $T^{3}$ term in the non-local radiation model is the key to reproduce this particular behavior. This pattern was reported in [19], where a Ricker function was proposed as an approximation for this dependency relationship. The proposed Ricker function accounts for a constant contribution that dominates at low densities, transitioning to an exponentially decaying function at higher densities. We note that the specific density value for this transition does not coincide with that in [19] because no calibration has been performed in our model. Differences between the three models considered herein are more noticeable in the region near the maximum ROS. For the ’only fuel’ model, the transition is very sharp, whereas the other models show a smoother transition.
@@
The rate of spread depends as well on the surface coverage, which is relevant for heterogeneous environments. Here, we model the surface coverage as a factor multiplying the fuel packing ratio, $R_{f,0}$, when assuming a theoretical maximum value $R_{f,max}$, as defined in in Eq. (4). The previous study was carried out for a wide range of fuel packing ratios. Here, we assume that the maximum packing ratio equals $R_{f,0}=0.01$ and analyze the propagation regime up to this value, where the transition from no fire to propagation conditions is located. Experiments show a nonlinear dependency of the rate of spread on the surface coverage [33]. Again, we compare the three different models described above for the dependency of the bulk density on the fuel load.
We compare three different models for the dependency of the bulk density on the fuel load. Eqs. (6) and (7) give the first approximation, where both, the gaseous and solid phase, are regarded. Additionally, their portion varies in time due to the combustion process. We name this modeling approach as ’mixture (time varying)’. The next approximation is given by Eq. (8), taking into account the mixture but not the time variation. As the simplest approximation, Eq. (9) does not depend on the gaseous phase but regards only the fuel (solid phase).
Figure 14: Rate of spread (ROS) depending on the surface coverage (SC) for three different fuel models. The simulations are without moisture $M=0$ and wind $w=0$.
@@
The key aspect when considering local radiation lies in the strong nonlinearity of the diffusion term. In [42], we already showed that a variation of the diffusion coefficient produces a change in the ROS. In the present model, the dependency of the diffusion mechanism on $T^{3}$ allows to connect the ROS with the energy contained in the medium, which is proportional to its temperature. This dependency is required to produce physically consistent results, reproducing the trends observed in all the cases discussed above.

### v2

@@ 3.2 Time stepping
The Strong Stability Preserving Runge–Kutta 3 (SSP-RK3) scheme [27, 52] for stepping in time the semi-discrete Eq. (72) is used. It is expressed as follows
@@ 4 Numerical results
One-dimensional (1D) cases are first considered to assess the influence of different parameters. For the 1D simulations presented herein, the computational domain is $[0,500]$ m and the initial conditions are
@@
For all cases, the computational mesh is composed of $N_{x}=1600$ cells and $\mathsf{CFL}=0.1$. Default parameters from Tabs. 24 are used, unless other configuration is specified.
@@ 4.1 Comparison of combustion functions
In Sec. 2.5, two approaches for the combustion function were presented. Using the fixed parameters in Tab. 4, the parameter $A_{c}$ of the point-wise constant combustion function in Eq. (42) is determined such that it approximates the non-linear Arrhenius law Eq. (40) in the relevant temperature domain. Fig. 2(a) shows this comparison. The relevant temperature domain is $[T_{pc},T_{\mathrm{up}}]$, where $T_{\mathrm{up}}$ approximates an upper bound for temperatures in a fire.
@@
Figure 2: Comparison of the simulations using the two combustion functions. (a) Comparison of the combustion function using the Arrhenius law in Eq. (40) and with a constant term in Eq. (42). (b) Solutions of the ADR model (32) using the two combustion functions. The graphs widely overlap.
For comparing the influence of the combustion function, the advection term is set zero ($w=0$) and the solutions of the ADR-model are computed for a moisture level $M=0.1$. The direct comparison of the solutions in Fig. 2(b) shows that both, the traveling wave profile and the traveling wave speed are almost identical for the two combustion functions. The maximal temperature is slightly higher for the combustion function using the Arrhenius law. This difference has its origin in the approximation in Fig. 2(a), where the constant combustion function has a lower value for large temperatures than the Arrhenius combustion function. The differences of the solutions can be reduced even further by applying calibration techniques for the parameter $A_{c}$. Because of the good approximation, the assumption of a constant reaction rate instead of the Arrhenius rate is convincing.
@@
In the further investigations, the piecewise constant reaction rate in Eq. (42) is used instead of the non-linear Arrhenius law.
@@ 4.2 Influence of radiation models
The radiation leads to a non-linear diffusion term, compare Eqs. (21) and (28). When studying the influence of various mechanisms on the system’s behavior in [45], the non-linear diffusion by radiation was not considered, and the diffusion was modeled only by conduction leading to a linear diffusion term with a modified larger constant $k_{c}$. Here, the differences in solutions caused by the simplification of a purely linear diffusion compared to the combined effect of radiation and conduction are demonstrated first.
Figure 3: Comparison of the solution profiles for the temperature $T$ and the fuel $Y$ at a fixed time with and without radiation. The parameter $k_{c}$ differs in the two cases for compensating the lack of radiation.
@@ 4.3 Influence of fuel moisture models
The different models for the fuel moisture in Sec. 2.6 are now compared for some specific scenarios.
@@ 4.3.1 Comparison of three moisture models
First, consider a green wood to total wood ratio $r=1$, meaning that all wood is green wood, and a moisture content of $M=0.1$ under no wind conditions. The rest of the parameters are set as default. In this case, the only difference between the simple and the complete two-stage moisture models S2M and C2M is that the latter considers a correction factor, $A_{i}$, in the specific heat approximation Eq. (66) below wood fibre saturation (moisture content below 0.3). Note that the value of $M=0.1$ used in this test case would be unrealistic for green wood, but it has been chosen as an example for comparison of the models.
@@
Figure 4: Comparison of the traveling wave profiles for the temperature $T$ and the fuel $Y$ at a fixed time for two-stage moisture models (S2M, C2M) with the three-stage moisture model S3M. The moisture content is $M=0.1$, the wind speed is $w=0$, and the live/dead ratio is $r=1$.
@@
The phase portraits of the solutions provide further insight into the traveling wave profiles. They highlight differences in the solution behavior very well, compare the use in [45]. To represent the phase portraits, we fix a certain point $x$ where the traveling wave will pass after some time. This point $x$ should not be too close to the maximum of the initial data for ensuring that the traveling wave front is already in its stable shape. For this point $x$, the solution values $T(x,t)$ and $Y(x,t)$ are plotted over each other, leading to the phase plot in Fig. 5(left). The phase plot shows that the traveling wave front is following nearly the same $T$-$Y$-curve, independent of whether a two-stage or a three-stage model was used. Additionally, the time-dependent solutions for this fixed point $x=350$ m is shown, giving the profiles in Fig. 5(right). This figure shows again the slightly slower rate of spread of the two-stage models. Note that Fig. 5(right) shows the time dependence; note that a faster traveling wave reaches the fixed spatial location in a shorter time.
Figure 5: Comparison of traveling wave profiles for the temperature $T$ and the fuel $Y$ for two-stage and three-stage moisture models with fixed $x=350$ m, and parameters $M=0.1$, $r=1$, $w=0$. Left: Phase portrait of the solutions at a fixed $x$; right: time-dependent solutions showing the traveling wave profiles.
The next step is to better visualize the differences between the 2- and 3-stage fuel moisture models. The 3-stage model explicitly represents the latent heating process into account, resulting in some plateau in Fig. 1(b). This difference is not visible in the phase portrait in Fig. 5, because it takes place while $Y=1$, following a straight line with increasing temperature. In contrast, Fig. 6 shows very well the differences of the time-dependent temperature curves at a fixed $x=350$ m during the heating processes. The parameters are fixed as in the previous figures. The solution using the S3M model shows the latent phase with a slightly increasing plateau, due to the width of the evaporation interval in Eq. (53). This evaporation plateau leads to a slowed heating process for temperatures around the evaporation temperature $T_{w}$. Contrarily, the S2M and C2M models show a monotonically increasing heating process without the evaporation plateau. In these models, the effect of latent heating is treated as a sensible heating and therefore evaporation cannot be explicitly represented.
Figure 6: Time dependent heating process for the two-stage models S2M, C2M, and the three-stage model S3M. The evaporation plateau leads to a slowed heating process for temperatures around the evaporation temperature $T_{w}$. Fixed parameters are $x=350$ m, $r=1$, $M=0.1$ and $w=0$.
@@
In the next step, the S2M model is fixed and the effect of varying moisture levels is studied.
@@ 4.3.2 Influence of moisture content
The following investigations use the simple two-stage moisture model S2M in Sec. 2.6.2. Further, advection effects are not regarded, so $w=0$. In comparison to the last section, the moisture content $M$ is now varied.
Fig. 7 shows the solution at $t=4000$ s for moisture contents $M=0$, $M=0.1$ and $M=0.2$. These moisture content values are representative of dead fuels with relative humidity levels below 90%. The traveling wave profiles in Fig. 7 show a decrease in the maximal temperature and in the traveling wave speed with increasing moisture content $M$. The surviving biomass fraction $Y$ is identical for all three values of $M$ even though the maximal temperature is different in all cases. In the complex system behavior, the larger maximal temperature values are compensated by the faster traveling wave speeds. Note that this inverse dependency between maximal temperature and traveling wave speed is due to the non-linear radiation term. In previous investigations [40], it was shown that a higher rate of spread is connected to a larger remaining biomass fraction $Y$. In this case now, the higher maximal temperature would lead to a smaller remaining biomass fraction, but the faster rate of spread compensates this effect.
Figure 7: Traveling wave solution profiles at a fixed time $t$ for the two-stage moisture model S2M for different moisture levels. Fixed values are $r=1$ and $w=0$. High moisture levels lead to a slower rate of spread and a lower maximal temperature.
@@
The preheating time is non-linear related to the moisture level. In particular, for this parameter setting, a moisture level of $M\geq 0.4$ stops the traveling wave behavior. The longer the preheating time is, the slower are the traveling waves because it needs longer to start the combustion process when the traveling wave front pushes forward.
@@
Figure 8: Comparison of different moisture levels in the S2M model with fixed $x=350$ m, $w=0$, $r=1$. Left: phase portrait of the solutions at a fixed $x$; right: time-depending solution profiles showing the traveling wave front. Higher moisture level lead to a slower ROS and a lower maximal temperature.
@@
Figure 9: Time-dependent temperature curves for three neighboring discretization cells (upper), and specific heat values in the three cells (lower). The moisture model is S2M with $M=0.2$, $w=0$, $r=1$ and $x=350$ m. The curves show a similar behavior with some time delay. The discontinuous values of $c_{p}$ lead to bumps in the temperature curves.
The occurrence of bumps is independent of the discretization parameters and desired for describing the temperature dependent evaporation process. They do not change the traveling wave behavior, see Fig. 8. Therefore, they are noticed as a feature of the model, not distracting any other observations.
@@ 4.4 Investigation of the parameter-dependency of ROS
The S2M model is used for some further investigations on the dependency of the rate of spread on the moisture content, the wind velocity, the relative humidity and the surface coverage. These investigations lead the path towards more realistic scenarios and deepen the understanding of the influence of moisture models.
@@
Fig. 10(a) shows the rate of spread (ROS) of the model in Eq. (32) with respect to the total moisture content, $M$, computed using the model S2M in Eq. (58). In the case without wind, $w=0$, the traveling wave stops for larger moisture contents. The fire front stops and the fire is not spreading any further. This solution’s behavior is different for situations with a small wind, $w=1$ or $w=2$ m/s. In these cases, the moisture content slows down the rate of spread but does not lead to a stopping of the fire front.
@@
Figure 10: Dependency of the rate of spread (ROS) on the moisture content $M$ and the wind velocity $w$. The S2M model is used for modeling the moisture. Higher moisture levels lead to smaller rate of spread while a higher wind velocity increases the rate of spread. If the wind is zero and the moisture level is high, the traveling wave stops.
@@
The observation that the traveling wave solution with some wind $w>0$ does not stop is in accordance with practical observations, although the underlying mechanisms differ in reality and in this model. In reality, a strong wind may lead to a heating of the fuel far ahead of the traveling wave front and therefore to an evaporation of the fuel moisture before the fire front arrives. This long-distance effect, however, is not represented in the current mathematical model. Here, the model only considers local radiative effects through the Rosseland approximation, which acts as a diffusion mechanism. As a result, heat is transferred between adjacent areas in a manner similar to conduction. To properly model this effect, the non-local effect of radiation should be considered as proposed in [25, 55]. This implies the resolution of additional equations for the radiation, thus increasing the complexity of the model.
@@
The impact of increasing fuel moisture on the ROS would be higher if, instead of considering a constant factor $\beta$, the expression in Eq. (39) is used, introducing a dependency $\beta=\beta(M)$. As $M$ increases, $\bar{c}_{p,f}$ also increases, leading to a decrease in $\beta$, which results in a larger reduction of the ROS for high moisture and wind values. Nevertheless, given the level of simplification in our model, we limit ourselves to a constant value for $\beta$ and defer this option for future research. This approach could improve the results for strong wind situations, where the effect of moisture is underestimated.
@@
In Sec. 2.3, the connection between wind velocity and topographies with constant slopes was drawn. Recall the key idea: Even though the investigations in Fig. 10 were carried out for a wind velocity, they are valid as well for a non-flat topography with a constant slope. Further effects of the topography on the spread of fire will be studied with simulation in two spatial dimensions in Sec. 4.5.
@@
In the next step, the relative humidity and the live to dead ratio $r$ of the fuel are varied. The live to dead ratio is relevant only for the complete two-stage moisture model C2M, which will be used for the following investigations.
@@
Figure 11: Dependencies of the rate of spread (ROS) with the C2M moisture model on the relative humidity RH ($\phi$) with varying wind velocity $w$ for $r=0.0$ (a), and varying green wood moisture content $M_{g}$ for $r=0.5$ (b). An increase of the relative humidity leads to a slower rate of spread in both variations.
@@
The two figures show a behavior that is expected from literature: In [38, Fig. 6], the comparison of different (empirical) models with simulation results is shown. All the models there show a decaying tendency. The different empirical models describe the decay either by a linear or an exponential decay function. The model results differ from this behavior in the case of very high relative humidity values, where an additional drop occurs. As a relative humidity near 100% is not the typical wildfire condition, the fit of the model to the literature values is still good.
@@
The dependence of the rate of spread on the bulk density of the fuel has been a widely studied topic for decades [16, 21]. This dependency is highly significant in heterogeneous environments, and recent ADR models have demonstrated the ability to accurately reproduce it [64]. Here, the focus is on studying the influence of the fuel bulk density, $\bar{\rho}_{f}R_{f,0}$ in Eq. (9), in the ROS. For this numerical experiment, a fixed fuel density is considered (see Tab. 2) and the fuel volume ratio, $R_{f,0}$, also called packing ratio in this context, is changed.
Three different models for the dependency of the bulk density on the fuel loadare compared. Eqs. (6) and (7) give the first approximation, where both, the gaseous and solid phase, are regarded. Additionally, their portion varies in time due to the combustion process. This modeling approach is referred to as “mixture (time varying)”. The next approximation is given by Eq. (8), taking into account the mixture but not the time variation. As the simplest approximation, Eq. (9) does not depend on the gaseous phase but regards only the fuel (solid phase).
Figure 12: Rate of spread (ROS) depending on the bulk density and the packing ratio for the three different models. The simulations are without moisture $M=0$ and wind $w=0$. For sufficiently large fuel bulk densities the rate of spread decreases with increasing fuel bulk densities.
Fig. 12 shows the dependency of the rate of spread on the fuel bulk density for the three approaches. The results evidence that there are two well-defined regions. For lower values of the fuel bulk density (and packing ratio), there is a transition from no fire conditions to wave propagation, increasing the ROS up to a maximum value. For larger densities (and packing ratio), the ROS decreases slowly, in accordance with the results reported in the literature. We believe that the role of the $T^{3}$ term in the non-local radiation model is the key to reproduce this particular behavior. This pattern was reported in [21], where a Ricker function was proposed as an approximation for this dependency relationship. The proposed Ricker function accounts for a constant contribution that dominates at low densities, transitioning to an exponentially decaying function at higher densities. The specific density value for this transition does not coincide with that in [21] because no calibration has been performed in our model. Differences between the three models considered herein are more noticeable in the region near the maximum ROS. For the ’only fuel’ model, the transition is very sharp, whereas the other models show a smoother transition.
@@
The rate of spread depends as well on the surface coverage, which is relevant for heterogeneous environments. Here, the surface coverage is modeled as a factor multiplying the fuel packing ratio, $R_{f,0}$, when assuming a theoretical maximum value $R_{f,max}$, as defined in in Eq. (4). The previous study was carried out for a wide range of fuel packing ratios. Here, the maximum packing ratio is assumed to be $R_{f,0}=0.01$, and the propagation regime up to this value is analyzed to identify the transition from no fire to propagation conditions. Within this range, the fuel load can be considered a limiting factor for propagation. Experiments show a non-linear dependency of the rate of spread on the surface coverage [36]. Again, the three different models described above for the dependency of the bulk density on the fuel load are compared.
Figure 13: Rate of spread (ROS) depending on the surface coverage (SC) for three different fuel models. The simulations are without moisture $M=0$ and wind $w=0$. The fire stops propagating for very small surface coverage.
@@
The key aspect when considering local radiation lies in the strong non-linearity of the diffusion term. In [45], it was shown that a variation of the diffusion coefficient produces a change in the ROS. In the present model, the dependency of the diffusion mechanism on $T^{3}$ allows to connect the ROS with the energy contained in the medium, which is proportional to its temperature. This dependency is required to produce physically consistent results, reproducing the trends observed in all the cases discussed above.